2{3x-1}=4 OR 2 square root of 3x-1 equals 4
Answer by fractalier(6550)

Okay from
2sqrt(3x - 1) = 4
we divide both sides by two first
sqrt(3x - 1) = 2
then square both sides
3x - 1 = 4
then solve and check
3x = 5
x = 5/3
It checks.